What current flows through a 2.54 cm diameter rod of pure silicon that is 20.0 cm
long, when 1.00 × 103
1
Expert's answer
2021-07-09T08:31:58-0400
"I=\frac{V}{R}\\I=\frac{\pi r^2V}{\rho l}\\
I=\frac{V}{R}\\I=\frac{\pi (0.0127)^2(1000)}{(2300) (0.2)}=1.1\ mA"